OpenStudy (anonymous):

A fair coin is tossed 3 times. (a) Find the probability that exactly 3 heads appear (b) Find the probability that no heads appear.

OpenStudy (anonymous):

Okay, so assuming two events are independent, meaning the outcome of one doesn't effect the outcome of the other, then the probability of both events happening is just the product of their probabilities. In other words if events A and B don't effect each other, then P(A and B) = P(A) * P(B) If you flip a coin twice, do the flips effect each other? Then what is the likely hood that you'll flip two heads?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Any thoughts? Work through this with me. I'm not just telling you the answers.

OpenStudy (anonymous):

a)1/2*1/2*1/2=1/8

OpenStudy (anonymous):

b)1/8

OpenStudy (anonymous):

Maybe Hashir will though, since he hates learning.

OpenStudy (anonymous):

@smoothmath i think it is a very simple question and it needs no great theoritical explanations............is casemac know basic theory she woly be able to understand it...and i m sure she is

OpenStudy (anonymous):

i understand a but i don't understand b

OpenStudy (anonymous):

If she understood how to do it, then she wouldn't have had to ask for help now, would she?

OpenStudy (anonymous):

ok look if no heads apper mean that every time we get a tail insted of head

OpenStudy (anonymous):

i thought for be it would be 10/2^3

OpenStudy (anonymous):

The first thing to understand about coin flips is that they are independent events. Flipping a head the first time doesn't make you more or less likely to flip heads the second time. So the probability of flipping three heads in a row is just P(heads) * P(heads) * P(heads) = .5 *.5 *.5 = .125 Similarly, the probability of flipping no heads just means the probability of flipping three tails in a row which is P(tails) * P(tails) * P(tails) = .5 *.5 *.5 = .125

OpenStudy (anonymous):

Why 10/2^3? And do you mean you thought it would be (10/2)^3?

OpenStudy (anonymous):

@smoothmath i guess u were right she dont know the basics........

OpenStudy (anonymous):

You're a noble man. I apologize if I was too aggressive.

OpenStudy (anonymous):

Casemac? How you doin'? Can we explain more? Have any questions?